# Instability of Legendrian knottedness, and non-regular Lagrangian concordances of knots

### v1

@@ 4.1. First strategy: Spinning the ambient Legendrian surgery discs
Since there is a canonical identification $(J^{1}\eta_{i}\times T^{*}S^{n},dz+p\,dq+\lambda_{S^{n}})=J^{1}(\eta_{i}\times S^{n}),dz+p\,dq+\lambda_{S^{n}})$ of contact manifolds, a product of a contact standard neighbourhood of $\eta_{i}$ with a Weinstein neighbourhood of $0_{S}^{n}$ is a standard neighbourhood of the Legendrian $\eta_{i}\times 0_{S^{n}}$. Any two contact neighbourhoods of $\eta_{i}\times 0_{S^{n}}$, $i=0,1,$ are contact isotopic; here we use the uniqueness of contact neighbourhood together with the assumption that these two Legendrians are isotopic.
@@
Now consider the Legendrian link $\Lambda\subset\mathbb{R}^{3}$ consisting of two unlinked standard unknots, with the surgery arc $\eta_{m}\subset\mathbb{R}^{3}$ as shown in Figure 17 or 18. In particular, the ambient Legendrian surgery produces the Legendrian knot $\Lambda_{\eta_{m}}=\Lambda_{m}$. Then consider the Legendrian surgery arc $\eta_{0}$ for the same Legendrian link $\Lambda$ depicted in Figure 19, for which $\Lambda_{\eta_{0}}$ clearly is Legendrian isotopic to the standard unknot $U$. Note that $\eta_{0}$ and $\eta_{m}$ both are stabilised and Legendrian isotopic in $\mathbb{R}^{3}$ relative boundary. However, any isotopy must clearly intersect $\Lambda$.
In the case when $n$ is odd, we can apply Corollary 4.3 directly to get the sought isotopy between the spun of $\eta_{m}$ and the spun of $\eta_{0}$. Lemma 4.5 then gives the sought Legendrian isotopy between the spuns.
In the case when $n$ is even, we can use the contact isotopy depicted in Figure 15 in order to verify the existence of a Legendrian isotopy that takes the arc $\eta_{m}$ to $\eta_{0}$, and for which the assumptions of Proposition 4.2 are met. Indeed, the obvious isotopy $\eta_{t}$ described in Figure 16 intersects $\Lambda$ twice, but with opposite signs of the corresponding double points on the induced trace cobordism. Again, corollary 4.3 finally shows the claim.
@@ 4.2. Second strategy: Spinning the pair of cancelling Weinstein handles
We apply Proposition 3.2 to the standard Lagrangian handle-attachment cobordism whose negative boundary is the two unlinked unknots $\Lambda$ and positive boundary is the Legendrian $\Lambda_{m}$, i.e. the result $\Lambda_{\eta_{m}}$ of an ambient Legendrian surgery on $\Lambda$ along the isotropic arc $\eta_{m}$. The result is a Weinstein structure on ×$\mathbb{R}\times\mathbb{R}^{3}$ consisting of a Weinstein one and two-handle in cancelling position for which the Lagrangian handle-attachment cobordism is tangent to the Liouville flow.
@@
In particular Proposition 3.2 gives us a Legendrian $\tilde{\Lambda}_{\eta}\subset\mathbb{R}^{3}_{sc}$ where the latter contact manifold is the result of a Weinstein 00-surgery on $\mathbb{R}^{3}$, such that $\tilde{\Lambda}_{\eta_{m}}$ becomes $\Lambda_{\eta_{m}}$ after a Weinstein 1-surgery on a knot $S\subset\mathbb{R}^{3}_{sc}\setminus\tilde{\Lambda}_{\eta_{m}}$ that cancels the 00-surgery. The attaching 00-sphere and cancelling $1$-sphere, as well as the knot $\tilde{\Lambda}_{\eta_{m}}$, are obtained from $\Lambda$ and $\eta_{m}$ by the recipe shown in Figure 14.
@@
There exists a stabilised Legendrian knot $S_{0}\subset(\mathbb{R}^{3})_{sc}\setminus\Lambda_{\eta_{m}}$ which also is in cancelling position, but for which $\Lambda_{\eta_{m}}$ becomes the standard unknot $U$ after cancelling the handles. For instance, we can use the ambient Legendrian surgery arc from Figure 19 and replace it with the corresponding attaching sphere from Proposition 3.2; again we can follow the recipe in Figure 14.
@@
Taking products with $T^{*}S^{n}$ results in a generalised Weinstein handle decomposition where $\mathbb{R}^{3}_{sc}\times T^{*}S^{n}$ are obtained from $\mathbb{R}^{3}\times T^{*}S^{n}$ by two subcritical surgeries along isotropic $n$-spheres, and where $\mathbb{R}^{3}\times T^{*}S^{n}$ is obtained from the latter by a generalised Weinstein surgery along the Legendrian $S\times 0_{S^{n}}$; this generalised surgery corresponds to attaching a generalised Weinstein handle of the form $T^{*}(D^{2}\times S^{n})$. Note that such a handle-attachment, as well as the corresponding contact surgery, is completely determined by a choice of Legendrian embedding of $D^{2}\times S^{n}$.
@@
Finally we can apply Corollary 4.4 to the two attaching tori $S_{0}\times 0_{S^{n}}$ and $S\times 0_{S^{n}}$ to deduce the claim that $\Lambda_{m}\times 0_{S^{n}}$ and $U\times 0_{S^{n}}$ are Legendrian isotopic.
@@
Alternatively, one can invoke the $h$-principle for flexible Lagrangian cobordisms [18, Theorem 4.2] to deduce that there is a symplectomorphism of the symplectisation of $\mathbb{R}^{3}\times T^{*}S^{n}$ that identifies the standard Lagrangian fillings of $U\times 0_{S^{n}}$ with the one of $\Lambda_{m}\times 0_{S^{n}}$ (given by the product of the one in Figure 2). The Legendrian isotopy is then constructed by applying Lemma 4.1 to the contactomorphism of the contact boundary that is induced by the symplectomorphism.
@@
Figure 19. Performing the Legendrian ambient surgery on two unlinked unknots, whose front projections moreover are disjoint, along an arc $\eta_{0}$ that intersects the two unknots precisely in their boundary points, produces the standard unknot $U$.
@@ 5. Constructing non-regular Lagrangian concordances
The existence of non-regular exact Lagrangian cobordisms were shown by Eliashberg and Murphy in [21] in high dimensions and Lin [31] in dimension four. Namely, these constructions produce exact Lagrangian cobordisms with an empty positive end inside a symplectisation; elementary topological considerations imply that such a cobordism cannot be regular. From this perspective, the example that we construct is different, it has a non-empty positive end and the simplest possible topology among Lagrangian cobordisms, namely that of a concordance.
First we need the following generalisation of [7, Theorem 3.2]
@@
Let $C\subset[-T,T]\times S^{3}$ be a smooth concordance that coincides with the cylinder over a knot $K_{\pm}$ near the boundary, and such that $[-T,T]\times S^{3}$ can be obtained by attaching $0,1,$ and $2$-handles onto a neighborhood of
($(\{-T\}\times S^{3})\cup C\subset[-T,T]\times S^{3}.$
@@
Consider the handle-decomposition of $[-T,T]\times S^{3}$ in which all handles are attached onto a neighbourhood ($(\{-T\}\times S^{3})\cup C$, and for which there are no $3$- or $4$-handles. We may also assume that there are no 00-handles, and that the $2$-handles are attached after the $1$-handles. Since $H_{*}([-T,T]\times S^{3},\{-T\}\times S^{3})=0$, it follows that the number of $1$-handles is equal to the number of $2$-handles; denote this number by $h$.
[…]
Attaching a four-handle to the end $\Sigma(K_{+})=S^{3}$ we obtain a cobordism $X$ from $\Sigma(K_{-})$ to $\emptyset$ which is simply connected. Indeed, turning it upside down, we have a cobordism with only 00-, $2$-, and $3$-handles.
[…]
The rest of the proof is the same as the argument in [7, Theorem 3.2]. In particular, this argument uses the adaptation [7, Theorem 3.1] of a result of Kronheimer and Mrowka that shows that that $\pi_{1}(\Sigma(K_{-}))$ admits a non-trivial representation $\rho$ in $SO(3)$ if and only if $K_{-}$ is not the unknot, while the argument in [29, Section I.3] shows that there can be no such representation under the assumption that $2h$ new generators and relations kill the group $\pi_{1}(\Sigma(K_{-}))$.
[…]
@@
It is unknown to the authors whether a regular Lagrangian cobordism is also decomposable. A decomposable Lagrangian cobordism $L\subset[-T,T]\times S^{3}$ is regular by a particular Weinstein handle decomposition of $[-T,T]\times S^{3}$ in which the handle decomposition of the neighbourhood $T^{*}L$ and the handles attached to $\{-T\}\times S^{3}\cup T^{*}L$ are in cancelling position. It is not clear if a general set of Weinstein handle attachment on $\{-T\}\times S^{3}\cup T^{*}L$ that produces $[-T,T]\times S^{3}$ can be deformed into such a position.
In recent work [8] by the first author it is shown that totally real concordances can be made Lagrangian after adding sufficiently many stabilisations to the Legendrian ends.
@@
Assume that there exists a totally real concordance from a Legendrian knot $\Lambda_{-}^{\prime}$ to $\Lambda_{+}^{\prime}$ (which according to Appendix A implies that $\Lambda_{\pm}^{\prime}$ are smoothly concordant, and with the same ${\tt tb}$ and ${\tt rot}$). Then this concordance can be deformed to a Lagrangian concordance from $\Lambda_{-}$ to $\Lambda_{+}$ by a smooth isotopy, where both $\Lambda_{\pm}$ are obtained from $\Lambda_{\pm}^{\prime}$ by performing a sufficiently large number $k\gg 0$ of stabilisations of both signs.
@@
If $C\subset\mathbb{R}_{t}\times\mathbb{R}^{3}$ is a decomposable Lagrangian concordance from $\Lambda_{-}$ to $\Lambda_{+}$ with $\tt{rot}(\Lambda_{\pm})=0$, then the reversed cobordism
@@
admits a compactly supported smooth isotopy to a concordance that is totally real for any choice of cylindrical almost complex structure.

### v2

@@ 4.2. First strategy: Spinning the ambient Legendrian surgery discs
Since there is a canonical identification
$(J^{1}\eta_{i}\times T^{*}S^{n},dz+p\,dq+\lambda_{S^{n}})=J^{1}(\eta_{i}\times S^{n}),dz+p\,dq+\lambda_{S^{n}})$
of contact manifolds, a product of a contact standard neighbourhood of $\eta_{i}$ with a Weinstein neighbourhood of $0_{S}^{n}$ is a standard neighbourhood of the Legendrian $\eta_{i}\times 0_{S^{n}}$. Any two contact neighbourhoods of $\eta_{i}\times 0_{S^{n}}$, $i=0,1,$ are contact isotopic; here we use the uniqueness of contact neighbourhood together with the assumption that these two Legendrians are isotopic.
@@
Now consider the Legendrian link $\Lambda\subset\mathbb{R}^{3}$ consisting of two unlinked standard unknots, with the surgery arc $\eta_{m}\subset\mathbb{R}^{3}$ as shown in Figure 17 or 18. In particular, the ambient Legendrian surgery produces the Legendrian knot $\Lambda_{\eta_{m}}=\Lambda_{m}$. Then consider the Legendrian surgery arc $\eta^{0}_{m}$ for the same Legendrian link $\Lambda$ depicted in Figure 19, for which $\Lambda_{\eta^{0}_{m}}$ clearly is Legendrian isotopic to the standard unknot $U$. Note that $\eta^{0}_{m}$ and $\eta_{m}$ both are stabilised and Legendrian isotopic in $\mathbb{R}^{3}$ relative boundary. However, any isotopy must clearly intersect $\Lambda$.
In the case when $n$ is odd, we can apply Corollary 4.2 directly to get the sought isotopy between the spun of $\eta_{m}$ and the spun of $\eta^{0}_{m}$. Lemma 4.4 then gives the sought Legendrian isotopy between the spuns.
In the case when $n$ is even, we can use the contact isotopy depicted in Figure 15 in order to verify the existence of a Legendrian isotopy that takes the arc $\eta_{m}$ to $\eta^{0}_{m}$, and for which the assumptions of Proposition 4.1 are met. Indeed, the obvious isotopy $\eta_{t}$ described in Figure 16 intersects $\Lambda$ twice, but with opposite signs of the corresponding double points on the induced trace cobordism. Again, Corollary 4.2 shows the claim.
@@ 4.3. Second strategy: Spinning the pair of cancelling Weinstein handles
We apply Proposition 3.2 to the standard Lagrangian handle-attachment cobordism whose negative boundary is the two unlinked unknots $\Lambda$ and positive boundary is the Legendrian $\Lambda_{m}$, i.e. the result $\Lambda_{\eta_{m}}$ of an ambient Legendrian surgery on $\Lambda$ along the isotropic arc $\eta_{m}$. The result is a Weinstein structure on $\mathbb{R}\times\mathbb{R}^{3}$ consisting of a Weinstein one and two-handle in cancelling position for which the Lagrangian handle-attachment cobordism is tangent to the Liouville flow.
@@
In particular Proposition 3.2 gives us a Legendrian $\tilde{\Lambda}_{\eta_{\color[rgb]{0,0,0}m}}\subset\left(\mathbb{R}^{3}\right)_{sc}$ where the latter contact manifold is the result of a Weinstein 0-surgery on $\mathbb{R}^{3}$, and such that $\tilde{\Lambda}_{\eta_{m}}$ becomes $\Lambda_{\eta_{m}}$ after a Weinstein 1-surgery on a knot $S\subset\left(\mathbb{R}^{3}\right)_{sc}\setminus\tilde{\Lambda}_{\eta_{m}}$ that cancels the 0-surgery. The attaching 0-sphere and cancelling $1$-sphere, as well as the knot $\tilde{\Lambda}_{\eta_{m}}$, are obtained from $\Lambda$ and $\eta_{m}$ by the recipe shown in Figure 13; we refer to Figure 18 for $\eta_{m}$.
@@
There exists a stabilised Legendrian knot $S_{0}\subset\left(\mathbb{R}^{3}\right)_{sc}\setminus\Lambda_{\eta_{m}}$ which also is in cancelling position, but for which $\Lambda_{\eta_{m}}$ becomes the standard unknot $U$ after cancelling the 1-handle with the 2-handle attached along $S_{0}$. This attaching sphere can be constructed analogously to $S$, but using the surgery arc $\eta^{0}_{m}$ shown in Figure 19 instead of $\eta_{m}$ shown in Figure 18. Here we again follow the recipe in Figure 13.
4.3.1. Digresson: Generalised Weinstein handle decompositions
The product of two Weinstein manifolds are naturally equipped with a handle-body structure where the handles of the product corresponds to products of handles. In particular, the number of handles for the induced handle-decomposition is the product of the number of handles of each factor. In the case when one of the factors is the cotangent bundle $T^{*}M$ of a closed smooth manifold there is a more economic decomposition, if one allows a generalisation of the Weinstein handle decomposition.
First recall that the standard Weinstein $k$-handle, as introduced by Weinstein in [41], in dimension $2n$ with $n\geq k$ is given by $T^{*}D^{k}\times\mathbb{C}^{n-k}$ with the standard symplectic structure. This handle can be attached to a Liouville domain by, roughly speaking, gluing the k1k-1-sphere
$\partial 0_{D^{k}}\times\{0\}\subset T^{*}D^{k}\times\mathbb{C}^{n-k}$
to a framed isotropic k1k-1 sphere in the boundary of a Liouville domain. The latter framed isotropic embedding is called the isotropic attaching sphere. A Weinstein manifold can be characterised as a Liouville domain that admits a handle-body decomposition consisting of standard Weinstein handles.
A generalised Weinstein $k$-handle is given by $T^{*}N^{k}\times\mathbb{C}^{n-k}$ where $D^{k}$ has been replaced by a general compact and smooth $k$-dimensional manifold $N^{k}$ with non-empty boundary. The generalised handles are similarly attached to a Liouville domain by gluing
$\partial 0_{N}\times\{0\}\subset T^{*}M^{k}\times\mathbb{C}^{n-k}$
along a framed isotropic embedding of $\partial N$ in the the contact boundary of the Liouville domain. We refer to the latter framed isotropic embedding as the isotropic attaching manifold; see [10] and [30] for more details on particular generalised handles. Those Liouville manifolds that admit a handle-body structure consisting of generalised Weinstein handles are obviously also Weinstien manifolds. By a generalised Weinstein handle body we mean a decomposition of a Weinstein manifold into generalised Weinstein handles.
A particular set of examples that admit natural generalised Weinstein handle decompositions is the product $X\times T^{*}M^{m}$ of a Weinstein manifold $X$ and a cotangent bundle $T^{*}M$ of a closed smooth $m$-dimensional manifold. Namely, this product has a simple decomposition into generalised Weinstein handles induced by any Weinstein handle decomposition of $X$. Namely, each Weinstein handle $T^{*}D^{k}\times\mathbb{C}^{n-k}$ in the handle-decomposition of $X$ gives rise to a generalised Weinstein handle $T^{*}(D^{k}\times M)\times\mathbb{C}^{n-k}$ in the generalised Weinstein handle-body decomposition of $X\times T^{*}M$. The topology of each handle in this case is thus $N^{k+m}=D^{k}\times M^{m}$.
4.3.2. A generalised Weinstein handle decomposition of the symplectisation $\mathbb{R}\times(\mathbb{R}^{3}\times T^{*}S^{n})$.
@@ 4.3.1. Digresson: Generalised Weinstein handle decompositions
In our situation we are interested in taking the product with $T^{*}S^{n}$. This results in a generalised Weinstein handle decomposition of the symplectisation $\mathbb{R}\times(\mathbb{R}^{3}\times T^{*}S^{n})$ obtained in the following manner. First, consider a Weinstein 1-handle attached on $\mathbb{R}^{3}$ and take the product with $T^{*}S^{n}$. This leads to a product Weinstein manifold with boundary $\left(\mathbb{R}^{3}\right)_{sc}\times T^{*}S^{n}.$ The latter contact manifold is obtained from $\mathbb{R}^{3}\times T^{*}S^{n}$ by a pair of subcritical Weinstein surgeries along isotropic $n$-spheres. Since the Weinstein 1-handle attachment that produces $\left(\mathbb{R}^{3}\right)_{sc}$ can be cancelled by a Weinstein 2-handle attached along $S\subset\left(\mathbb{R}^{3}\right)_{sc}$, taking a product of the latter handle and $T^{*}S^{n}$ then yields $\mathbb{R}^{3}\times T^{*}S^{n}$ obtained from $\left(\mathbb{R}^{3}\right)_{sc}\times T^{*}S^{n}$ by a generalised Weinstein surgery along the Legendrian $S\times 0_{S^{n}}$. Note that the latter generalised surgery corresponds to attaching a generalised Weinstein handle of the form $T^{*}(D^{2}\times S^{n})$. Note that such a handle-attachment, as well as the corresponding contact surgery, is completely determined by a choice of Legendrian embedding of $D^{2}\times S^{n}$.
@@
Finally we can apply Corollary 4.3 to the two attaching manifolds $S_{0}\times 0_{S^{n}}$ and $S\times 0_{S^{n}}$ to deduce the claim that $\Lambda_{m}\times 0_{S^{n}}$ and $U\times 0_{S^{n}}$ are Legendrian isotopic.
@@
Alternatively, one can invoke the $h$-principle for flexible Lagrangian cobordisms [19, Theorem 4.2] to deduce that there is a symplectomorphism of the symplectisation of $\mathbb{R}^{3}\times T^{*}S^{n}$ that identifies the standard Lagrangian fillings of $U\times 0_{S^{n}}$ with the one of $\Lambda_{m}\times 0_{S^{n}}$ (given by the product of the one in Figure 2). The Legendrian isotopy is then constructed by applying the below Lemma 4.5 to the contactomorphism of the contact boundary that is induced by the symplectomorphism.
Lemma 4.5.
Proof.
$\Phi_{t}\in\mathrm{Cont}(Y\times T^{*}S^{n}),$
$\Phi_{t}(y,\theta,p)=(\phi_{t}(y),\theta,p\cdot e^{g_{t}(y)})$
@@
Figure 19. Performing the Legendrian ambient surgery on two unlinked unknots, whose front projections moreover are disjoint, along an arc $\eta^{0}_{m}$ that intersects the two unknots precisely in their boundary points, produces the standard unknot $U$.
@@ 5. Constructing non-regular Lagrangian concordances
The existence of non-regular exact Lagrangian cobordisms were shown by Eliashberg and Murphy in [22] in high dimensions and Lin [33] in dimension four. Namely, these constructions produce exact Lagrangian cobordisms with an empty positive end inside a symplectisation; elementary topological considerations imply that such a cobordism cannot be regular. From this perspective, the examples that we construct here are different, since they have a non-empty positive end and the simplest possible topology among Lagrangian cobordisms; they are concordances.
First we need the following generalisation of [7, Theorem 3.2].
@@
Let $C\subset[-T,T]\times S^{3}$ be a smooth concordance that coincides with the cylinder over a knot $K_{\pm}$ near the boundary, and such that $T,T]\times S^{3}$ can be obtained by attaching $0,1,$ and $2$-handles onto a neighborhood of
$(\{-T\}\times S^{3})\cup C\subset[-T,T]\times S^{3}.$
@@
Consider the handle-decomposition of $T,T]\times S^{3}$ in which all handles are attached onto a neighbourhood $(\{-T\}\times S^{3})\cup C$, and for which there are no $3$- or $4$-handles. We may also assume that there are no 0-handles, and that the $2$-handles are attached after the $1$-handles. Since $H_{*}([-T,T]\times S^{3},\{-T\}\times S^{3})=0$, it follows that the number of $1$-handles is equal to the number of $2$-handles; denote this number by $h$.
Then we consider the double cover of $T,T]\times S^{3}$ branched over the concordance $C$. This branched cover thus has a handle decomposition for which there are $2h$ handles in the cover of index one as well as of index two. This branched double cover is a cobordism from $\Sigma(K_{-})$ to $\Sigma(K_{+})$, where $\Sigma(K_{\pm})$ denotes the double cover of $S^{3}$ branched along the knot $K_{\pm}$. Note that $\Sigma(K_{+})=S^{3}$ since $K_{+}$ is the unknot.
Attaching a four-handle to the end $\Sigma(K_{+})=S^{3}$ we obtain a cobordism $X$ from $\Sigma(K_{-})$ to $\emptyset$ which is simply connected. Indeed, turning it upside down, we have a cobordism with only 0-, 22-, and $3$-handles.
If we then consider $X$ as obtained by attaching $1$, 22, and $4$-handles on $\Sigma(K_{-})$, this means that the trivial group can obtained from $\pi_{1}(\Sigma(K_{-}))$ by adding $2h$ new generators (coming from the one-handles) and $2h$ new relations (coming from the two-handles).
The rest of the proof is the same as the argument in [7, Theorem 3.2]. In particular, this argument uses the adaptation [7, Theorem 3.1] of a result of Kronheimer and Mrowka that shows that that $\pi_{1}(\Sigma(K_{-}))$ admits a non-trivial representation $\rho$ in $SO(3)$ if and only if $K_{-}$ is not the unknot, while the argument in [31, Section I.3] shows that there can be no such representation under the assumption that $2h$ new generators and relations kill the group $\pi_{1}(\Sigma(K_{-}))$.
We finish by recalling the latter argument. The group $G_{0}\coloneqq\pi_{1}(\Sigma(K_{-}))/\ker\rho$ is a non-trivial group by the assumption that $\rho$ is non-trivial. This quotient group can be taken to have the same generators as $\pi_{1}(\Sigma(K_{-}))$ (in particular it is finitely generated). In addition, it admits a faithful non-trivial representation into $SO(3)$. Adjoining $2h$ new generators, i.e. considering the free product $G_{0}\star\mathbb{F}_{2h}$ of $G_{0}$ and the free group on $2h$ generators, together with the $2h$ relations as above, we obtain a group $G$. This group is the quotient of the free product $\pi_{1}(\Sigma(K_{-}))\star\mathbb{F}_{2h}$ by the normal subgroup generated by the $2h$ relations together with the relations from $\ker\rho$. In particular, $G$ is the quotient of the trivial group $\pi_{1}(X)$ and, thus, $G$ is itself trivial. Finally we can use [26, Theorem 3] to show that $G_{0}$ embeds in $G$, which means that $G_{0}$ is trivial, and thus leading to the sought contradiction. (Note that the condition on the determinant of the $2h$ relations needed in the cited theorem is satisfied here by the argument above, by which $H_{1}(\Sigma(K_{-}))$ is killed inside $H_{1}(X)$ by these $2h$ relations in the abelianisation.)
@@
It is unknown to the authors whether a regular Lagrangian cobordism is also decomposable. A decomposable Lagrangian cobordism $L\subset[-T,T]\times S^{3}$ is regular by a Weinstein handle decomposition of $T,T]\times S^{3}$ of a very particular type, in which the handle decomposition of the neighbourhood $T^{*}L$ and the handles attached to $\{-T\}\times S^{3}\cup T^{*}L$ are in cancelling position. It is not clear if a general set of Weinstein handle attachment on $\{-T\}\times S^{3}\cup T^{*}L$ that produces $T,T]\times S^{3}$ can be deformed into such a position.
In recent work [8] by the first author it is shown that totally real concordances can be deformed in order to become Lagrangian after adding sufficiently many stabilisations to the Legendrian ends. To formulate the result we need to recall the following definition. An almost complex structure on the symplectisation $(\mathbb{R}_{t}\times Y,d(e^{t}\,\alpha))$ of a contact manifold $(Y,\alpha)$ is called cylindrical if
  • $J$ is invariant under translation of the $t$-coordinate;
  • $J$ preserves the contact planes $\ker\alpha\subset TY$ in each level-set $\{t\}\times Y$ and is compatible with $d\alpha$ inside the same, i.e. $d\alpha(\cdot,J\cdot)$ defines a Riemannian metric on $\ker\alpha$; and
  • $J\partial_{t}=R_{\alpha}$ where $R_{\alpha}$ is the Reeb vector field.
@@
Assume that for a given choice of cylindrical almost complex structure on $(\mathbb{R}\times S^{3},{\color[rgb]{0,0,0}d(e^{t}\,\alpha_{st})})$ there exists a totally real concordance from a Legendrian knot $\Lambda_{-}^{\prime}$ to $\Lambda_{+}^{\prime}$ (which according to Appendix A implies that $\Lambda_{\pm}^{\prime}$ are smoothly concordant, and with the same $\tt tb$} and $\tt rot$}). Then this concordance can be deformed to a Lagrangian concordance from $\Lambda_{-}$ to $\Lambda_{+}$ by a smooth isotopy, where both $\Lambda_{\pm}$ are obtained from $\Lambda_{\pm}^{\prime}$ by performing a sufficiently large number $k\gg 0$ of stabilisations of both signs.
@@
For any decomposable Lagrangian concordance
$C\subset(\mathbb{R}_{t}\times\color[rgb]{0,0,0}\mathbb{R}^{3}_{xyz},d(e^{t}(dz-y\,dx)))$
from a Legendrian knot $\Lambda_{-}$ to a Legendrian knot $\Lambda_{+}$, the reversed cobordism
@@
admits a compactly supported smooth isotopy to a cobordism that is totally real, for any given choice of cylindrical almost complex structure.